Willie Company sells 32,000 units at $22 per unit. Variable costs are $13.86 per unit, and fixed costs are $153,700.

Determine (a) the contribution margin ratio, (b) the unit contribution margin, and (c) income from operations.

a. Contribution margin ratio (Enter as a whole number.) %
b. Unit contribution margin (Round to the nearest cent.) $ per unit
c. Income from operations $
Answer #1

Contribution margin=Sales-Variable cost

=(22-13.86)=$8.14 per unit

Contribution margin ratio=Contribution margin/Sales

=8.14/22=37%

Total Contribution margin=(8.14*32,000)=$260480

Less:Fixed costs=(153,700)

Income from operations=$106780

Contribution margin ratio 37%
Unit contribution margin $8.14 per unit
Income from operations $106780